OpenStudy (luigi0210):

A ladder 25 feet long is leaning against a vertical wall. The bottom of the ladder is being pulled horizontally away from the wall at 3 ft/sec. How fast is the angle between the ladder and the ground changing when the top of the ladder is 20 ft from the ground?

OpenStudy (goformit100):



OpenStudy (anonymous):



OpenStudy (anonymous):

\(b = 3t\\\cos\theta={b\over l}\)

OpenStudy (anonymous):

\[\cos\theta={3\over l}t\] differentiate with respect to "t" and you'd get \(\Large {d\theta\over dt}\)